OpenStudy (anonymous):

What is the equation of the line that passes through the points (-3, 2) and (-5, 8)? 3x - y = -11 3x + y = - 7 3x + y = 5 x + 3y = 3

OpenStudy (anonymous):

3x + y = -7

OpenStudy (mathmale):

BSandStorm: solution of this problem takes several steps. First, find the slope of the line that passes through the 2 given points. Then choose one of the points and substitute its x- and y-coordinates, along with the slope, m, into the point-slope form of the equation of a straight line: \[y-y _{0}=m(x-x _{0}).\] I strongly suggest that you do this yourself so as to understand the process. Good luck!

OpenStudy (anonymous):

i couldn't remember how to write out the process and didn't feel like shuffling notes from the Lord knows when, so I just answered it the TI-83 way. Answer the equation for y, graph, and look at the table. If the points match up, it's the right equation.

OpenStudy (mathmale):

The choice is yours. We could go through the process if you like, or you could wait 'til you feel more like reviewing and practicing this.